You plan to do a proliferation assay using 10 ml of cells that you prepared in A3. You decided to use Concavalin A mitogen to measure the proliferation of T cells using a Flow cytometry proliferation assay. Your stock of Con A stock solution is 100 micrograms/ml. The recommendation is to treat the cells with a concentration of 10ng/ml. What volume of Con A will you add from the stock solution into you 10 ml cells?
Answer #1

stock concentration (M1) = 100 micrograms/mL

concentration needed (M2)= 10 nanogram/mL

final volume (V2)= 10 mL

volume of CoA needed (V1)=?

1 nanogram=10^-3 microgram

M2= 10 \times 10^-3 microgram/mL= 10^-2 microgram/mL

M1V1=M2V2

V1= M2V2/M1

= (10^-2\times10)/100

= 10^-3 mL

1 mL= 1000 microliter

10^-3 mL= 10^3\times1000=1 microliter.

volume of Con A will you add from the stock solution into you 10 ml cells= 1 microliter.